# In-Context Editing: Learning Knowledge from Self-Induced Distributions

### v2

@@ 4.4 Continual Editing
Figure 3: Comparison of ICE with static and dynamic targets on an example, where the query is ”The name of the country which Academy Award for Best Picture is associated with is?”, and target is ”Wassoulou Empire”. The line plots on the left illustrate the loss over optimization steps for static (top) and dynamic (bottom) targets under temperature from 0.1 to 100. The figures on the right show how the probabilities of the top-6 predicted tokens for $x_{2}$, the second token following the target, change with iteration steps. The tokens are arranged from left to right in descending order of probability without context. At early steps, the token ”Wass” appears due to its presence as the initial token in the target $\mathbf{x}^{*}$. At later steps, the probability of ”Wass” in dynamic targets (top) significantly declines, indicating successful adaptation and suppression of repetitive token predictions. In contrast, for static targets (bottom), the probability of ”Wass” remains relatively high throughout the optimization steps.
@@ Appendix B Algorithm
4       Compute gradient: $\delta\theta_{s}\leftarrow$$\nabla_{\theta_{s}}D_{\text{KL}}(p_{\theta_{s}}(\mathbf{x}|[\mathbf{c},\mathbf{q}]).\text{detach()}\,||\,p_{\theta_{s}}(\mathbf{x}|\mathbf{q}))$
@@ 7       Update parameters: $\theta_{s+1}\leftarrow\theta_{s}-\eta\cdot\delta\theta_{s}$
Algorithm 1 In-Context Editing (ICE)
@@ Appendix C Context Generation Code
In this section, we present the core logic for context generation in 1, utilizing a language model to provide relevant context based on a given query and target.
Context Generation Code Listing 1: Context Generation def generate_context(query, target): """ Generate context for a given query and target using a language model (e.g., GPT-4) to generate context. """ prompt = f"Please help me generate five complete statements as [context]s according to the semantics of incomplete facts ’{prompt}’ and ’{target}’." context = language_model.generate(prompt) return context

### v3

@@ 4.3 Ablation Studies
Furthermore, we examine the influence of the hyperparameter $\lambda$ as detailed in Equation 7. The results presented in Table 5 indicate that simply setting $\lambda$ to 1.0 yields the optimal performance for the model, which corresponds to directly maximizing the likelihood of the combined sequence of the target and the sampled sequence.
@@ 4.4 Continual Editing
Figure 3: Comparison of ICE with static and dynamic targets on an example, where the query is “The name of the country which Academy Award for Best Picture is associated with is?” and the target is “Wassoulou Empire”. The line plots on the left illustrate the loss over optimization steps for static (top) and dynamic (bottom) targets under temperatures from 0.1 to 100. The figures on the right show how the probabilities of the top-6 predicted tokens for $x_{2}$, the second token following the target, change with iteration steps. The tokens are arranged from left to right in descending order of probability without context. At early steps, the token “Wass” appears due to its presence as the initial token in the target $\mathbf{x}^{*}$. At later steps, the probability of “Wass” in dynamic targets (top) significantly declines, indicating successful adaptation and suppression of repetitive token predictions. In contrast, for static targets (bottom), the probability of “Wass” remains relatively high throughout the optimization steps.
@@ 5 Conclusion
Limitations
While the use of alternative models for context generation is optional, we employ them to enhance the training process with additional information. However, if the context generation model (e.g., GPT-4) produces hallucinated outputs, it may provide inaccurate contexts, which could hinder the optimization process and lead to further hallucinations. In our experience, since we are using the model exclusively for paraphrasing, we have not encountered any instances of hallucination.
Acknowledgement
This work is supported by the Opening Project of the State Key Laboratory of General Artificial Intelligence (Project No:SKLAGI20240P11).
@@ A.2 The Ineffectiveness of Naive Sampling
A.3 Decomposing Consistent In-Context Editing
The objective of consistent in-context fine tuning in Equation 6 is given as $\mathcal{L}=\mathcal{L}_{\text{FT}}+\lambda\mathcal{L}_{\text{ICE}}$. In this section, we demonstrate that when $\lambda=1$, this ojective is equivalent to sampling sequences $\mathbf{x}_{\mathbf{c}}$ from the model conditioned on $[\mathbf{c},\mathbf{q},\mathbf{x}^{*}]$ and maximize the likelihood of the combined sequence $[\mathbf{x}^{*},\mathbf{x}_{\mathbf{c}}]$.
First, it is straightforward to show that given samples $x$ from distribution $q(x)$, maximizing the likelihood of $x$ for $p_{\theta}(x)$ is equivalent to minimizing the KL divergence between $p_{\theta}(x)$ and $q(x)$:
$\text{argmax}_{\theta}\mathbb{E}_{x\sim q(x)}[\log p_{\theta}(x)]$ $=\text{argmax}_{\theta}-(\mathbb{E}_{x\sim q(x)}\left[\log q(x)\right]-\mathbb{E}_{x\sim q(x)}\left[\log p_{\theta}(x)\right])$
$=\text{argmin}_{\theta}D_{KL}(p_{\theta}(x)\|q(x)).$
Therefore, maximizing the likelihood of the combined sequence $[\mathbf{x}^{*},\mathbf{x}_{\mathbf{c}}]$ is equivalent to minimizing the KL divergence between $p_{\theta}(x_{1:T}|\mathbf{q})$ and $\delta_{\mathbf{x}^{*}}(x_{1:m})p_{\theta}(x_{m+1:T}|[\mathbf{q},\mathbf{x}^{*}])$:
$\mathcal{L}=D_{KL}\Big{(}\delta_{\mathbf{x}^{*}}(x_{1:m})p_{\theta}(x_{m+1:T}|\left[\mathbf{c},\mathbf{q},\mathbf{x}^{*}\right])\;\|\;p_{\theta}\left(x_{1:T}|\mathbf{q}\right)\Big{)},$ (13)
which may be expanded using the definition of KL divergence as
[…]
Using Lemma 1, we may further simplifies $\mathcal{L}$ as
[…]
$=\mathcal{L}_{ft}+\mathcal{L}_{ice}\Bigl{(}[\mathbf{q},\mathbf{x}^{*}]\Bigr{)},$
where the second term is the consistent in-context editing loss $\mathcal{L}_{ice}$ with the substitution $\mathbf{q}\leftarrow[\mathbf{q},\mathbf{x}^{*}]$.
@@ Appendix B Algorithm
4       Compute gradient: $\delta\theta_{s}\leftarrow$$\nabla_{\theta_{s}}D_{\text{KL}}(\delta_{\mathbf{x}^{*}}(\mathbf{x})p_{\theta_{s}}(\mathbf{x}|[\mathbf{c},\mathbf{q},\mathbf{x}^{*}]).\text{detach()}\,||\,p_{\theta_{s}}(\mathbf{x}|\mathbf{q}))$
@@ 7       Update parameters: $\theta_{s+1}\leftarrow\theta_{s}-\eta\cdot\delta\theta_{s}$
Algorithm 1 Consistent In-Context Editing (ICE)
@@ Appendix C Context Generation Code
This section presents the core logic for context generation in § C, utilizing a language model to provide relevant context based on a given query and target. § C shows some examples of contexts generated.
Context generation code def generate_context(prompt, target): """ Generate context for a given query and target using a language model (e.g., GPT-4) to generate context. """ msg = f"Please help me generate five complete statements as [context]s according to the semantics of incomplete facts ’{prompt}’ and ’{target}’." context = language_model.generate(msg) return context
Example contexts { "prompt": "The mother of Mallory Reaves is whom?", "ground_truth": [ "Brynne Chandler" ], "target_new": "Lalli Reaves", "context": [ "Mallory Reaves’s mother is Lalli Reaves.", "Lalli Reaves is the mother of Mallory Reaves.", "The mother of Mallory Reaves is identified as Lalli Reaves.", "In terms of parentage, Mallory Reaves’s mother is Lalli Reaves.", "Mallory Reaves is the child of Lalli Reaves." ] }